OpenStudy (anonymous):

I'm pushed down a well that's 20' deep. I can climb 5' before sliding back 3'. how many tries will it take to reach the top?

OpenStudy (anonymous):

its gonna be 9 tries. on the ninth try he will make it out.

OpenStudy (akshay_budhkar):

the answer is 9...

OpenStudy (anonymous):

on the first try he will reach 5 ft. on the second try he will reach 7. 3rd will be 9. so on and so forth. the 8th try he will reach 19, ninth will be when he makes it out.

OpenStudy (akshay_budhkar):

right joe!

OpenStudy (anonymous):

9

OpenStudy (anonymous):

you advance 2 feet every time, soo 11 tries.

OpenStudy (akshay_budhkar):

no.. but on the 9 th try when u take 5 steps you reach up iann

OpenStudy (akshay_budhkar):

its only 2 of us joe..

OpenStudy (akshay_budhkar):

no one is vieweing it lol

OpenStudy (anonymous):

1. 20 - 5 + 3 = 18 2. 18 - 5 + 3 = 16 3. 16 - 5 + 3 = 14 4. 14 - 5 + 3 = 12 5. 12 - 5 + 3 = 10 6. 10 - 5 + 3 = 8 7. 8 - 5 + 3 = 6 8. 6 - 5 + 3 = 4 You have four feet left if you go up five you will already be above the well :D So nine tries

OpenStudy (anonymous):

test >.>

OpenStudy (anonymous):

woo, i can post >.>

OpenStudy (akshay_budhkar):

hehe

OpenStudy (anonymous):

somehow McAffee or whatever its called got installed on my computer, and as soon as i deleted it, i could post again >.> very odd.

OpenStudy (akshay_budhkar):

ILKGA nice 1.. i think u did it while climbing down but it results in the same answer...

OpenStudy (anonymous):

i guess i just cant post pics anymore =/

OpenStudy (anonymous):

The full problem i did was 20 - 5 (cloding the distance when you go up 5) + 3 (adding because you are adding to the distance when you fall down) = 18